TanA × Tan2A = 1

Prove sin3A + cos3A = 1
1
Expert's answer
2020-11-29T19:26:36-0500

"\tan x * \tan 2x = 1\\
\dfrac{\sin x}{\cos x} * \dfrac{\sin 2x}{\cos 2x} = 1\\ \ \\
\dfrac{\sin x}{\cos x} * \dfrac{2\sin x *cosx}{\cos^2 x-\sin^2x} = 1\\ \ \\
\ \dfrac{2\sin^2 x}{\cos^2 x-\sin^2x} = 1\\
2\sin^2 x = \cos^2 x - \sin^2 x\\
therefore \ \cos^2 x = 3\sin^2 x\\
\cos x = \pm \sqrt3\sin x\\
\ \\
\ \\
\sin 3x +\cos 3x = \\
= \sin x*\cos 2x +\sin 2x * \cos x + \\
\cos x *\cos 2x -\sin x *sin2x = \\
= \sin x *\cos^2x -\sin^3x + 2\sin x*\cos^2x + \\
+\cos^3 x -\sin x*\cos^2x - 2 \sin^2x* \cos x = \\
 = \cos^3x-\sin^3x+2\sin x\cos^2x-2\sin^2 x\cos x = \\
= (\cos x-\sin x)(\cos^2x+\sin x\cos x +\sin^2 x) +\\
+2\sin x\cos x(\cos x-\sin x) = (\cos x-\sin x)(1+ 3\sin x\cos x) = \\
= \cos x-\sin x +3\sin x\cos^2 x -3\sin^2 x\cos x = \\
1) \sqrt3\sin x -\sin x + 3\sin^3 x -3\sqrt3\sin^3x = \\
= -\sqrt3\sin x(3\sin^2 x -1) +\sin x(3\sin^2x-1) =\\
(3\sin^2x-1)(\sin x-\sqrt3\sin x) \ne1 \\
2) -\sqrt3\sin x -\sin x + 3\sin^3 x +3\sqrt3\sin^3x = \\
= \sqrt3\sin x(3\sin^2 x -1) +\sin x(3\sin^2x-1) =\\
(3\sin^2x-1)(\sin x+\sqrt3\sin x) \ne1 
\\
\ \\
answer: \sin3x+\cos3x \ne 1"